# Decimetric Type-IV Radio Bursts of Solar Flares Induced by Electron Cyclotron Maser Emission

### v1

@@ Abstract
The radiation mechanism of decimetric wideband and pulsating radio bursts from the Sun (in terms of decimetric type-IV (t-IVdm) burst) and other flaring stars is a long-standing problem. Early investigations were based on the leading-spot hypothesis for the sun and yielded contradictory results. Here we provide essential evidence to support the scenario that these bursts are induced by the electron cyclotron maser emission (ECME) driven by energetic electrons mirrored in converging sunspot fields. This is done by analyzing 2 flare-associated t-IVdm bursts that are selected from a larger sample of 60 bursts recorded during 2010–2014, according to the levels of polarization and whether the source-field orientation can be unambiguously determined. We further modeled the transport of downward-streaming energetic electrons along a coronal loop and found most electrons get mirrored within the specific altitude range of 20–100 Mm. This explains why such bursts tend to have well-defined spectral ranges. The study greatly expands the application of ECME in solar radio astronomy and provides solar samples for similar bursts from other flaring stars.
@@ 1 Introduction
Due to the close proximity of the Sun to the Earth, solar radio astronomy has the advantage of high spectral and spatial observations. This favors the determination of the underlying radiation mechanism. Even so, the exact origin of solar t-IVdm bursts and their substantial fine structures remain largely unresolved (Bernold, 1980; Benz, 1980; Slottje, 1982; Dulk, 1985; Güdel & Benz, 1988; Isliker & Benz, 1994).
To determine the primary radiation mechanism, it is essential to identify the nature of the emission mode in terms of two freely escaping magnetoionic modes: extraordinary (X) or ordinary (O). This requires knowledge of the sense of polarization and orientation of the source magnetic field. Due to the lack of imaging data at decimetric wavelengths, to determine the source field orientation earlier studies (Aschwanden, 1986; Zlobec et al., 1987) were based on the leading sunspot hypothesis that assumes the bursts originate from the leading spots of the associated flares. Contradictory conclusions were drawn by these studies. As a result, distinct mechanisms have been proposed: plasma emission (PE; Kuijpers, 1974, 1980; Zaitsev & Stepanov, 1975; Benz & Kuijpers, 1976), electron cyclotron maser emission (ECME; Benz & Tarnstrom, 1976; Aschwanden & Benz, 1988a, b), or gyrosynchrotron radiation.
Over the past eight decades of solar radio astronomy, only one type of solar radio bursts — spikes — is convincingly attributed to ECME, while several other types of solar bursts such as type-I, II, and III have been attributed to PE. Recent studies have reported a rare ECME-like event with week-long decimetric pulsations. The event is not directly associated with any specific flare (Yu et al., 2024; Sharma et al., 2024). Here we investigated flare-induced t-IVdm bursts using extreme ultraviolet (EUV) and radio imaging observations. Among the 60 t-IVdm events recorded during 2010–2014, we found 2 flare-associated bursts with medium-strong levels of polarization and from sources near a sunspot where the field orientation can be determined. For both bursts, the maximum brightness temperature ($T_{B}$) exceeds 1010 K and their sources coincide with the sites of the second harmonic of the extrapolated local electron gyrofrequency. These findings provide substantial evidence that the flare-induced t-IVdm bursts are mainly the harmonic X-mode radiated via ECME. By modeling the electron transport along a given loop, we further suggest these bursts are excited by energetic electrons mirrored in converging sunspot fields and elaborate why such t-IVdm bursts tend to have well-defined spectral cutoffs. In Section 2, we introduce the data and event selection, followed by modeling methods in Section 3. The results are presented in Section 4 and the discussion is provided in Section 5.
@@ 2.1 Data and Instruments
The t-IVdm bursts were identified using the Bleien 7m radio telescope, which covers 175 to 870 MHz with a temporal resolution of 0.25 seconds. The Nançay Radioheliograph (NRH; Kerdraon & Delouis, 1997) offers imaging data with polarization measurements at 10 frequencies spanning from 150 to 445 MHz. The spatial resolution depends on the frequency and time of observation, being $\sim$2 at 445 MHz and $\sim$6 at 150 MHz in the summer, and up to three times larger along the north-south direction in the winter.
@@ 2.2 Event Selection
According to the general definition of t-IVdm bursts, we selected events with the maximum frequency exceeding 300 MHz, the bandwidth wider than 100 MHz, and the duration longer than 5 minutes. This yielded 60 events, from which we selected 37 flare-associated events (Table 1) according to the following criteria: the burst should be within 30 minutes around the peak of a flare, which originates from the active region associated with the NRH sources.
To identify the radiation mode, we only retained events with a high to moderate level of polarization ($>$30%) whose sources are located near a sunspot, with both latitude and longitude within 60, to ensure an unambiguous determination of the source-field orientation. The field orientation is further verified by the nonlinear force-free field (NLFFF) extrapolation. After applying these criteria, we identified 2 events.
3 Modeling Methods
@@ 3.1 Magnetic Field Processing and Extrapolation
We analyzed the HMI vector magnetic field data (“hmi.B_720s”), which cover the full solar disk with a temporal cadence of 720 s and a spatial resolution of 0.505′′ (Schou et al., 2012). The coronal magnetic field was reconstructed using the magneto-frictional method (Guo et al., 2016). To do this, we first corrected the projection effect (Gary & Hagyard, 1990) since the active region is away from the disk center, and then removed the residual force on the magnetograms using a preprocessing method (Wiegelmann et al., 2006). We applied two metrics (Wheatland et al., 2000) to check whether the magnetic field reaches the force-free state or not. The obtained force-freeness metrics for both events are lower than 0.25, and the divergence-freeness metrics are close to 8 $\times$ 10-5 , similar to previous studies (Zhong et al., 2019).
3.2 Particle Transport Modeling
@@ 4.1 Multiwavelength data analysis for the 20110924 event
Fig. 1 presents the spectra of the t-IVdm burst, along with the NRH data for $T_{B}$ and polarization. The event occurred from 12:50 to 13:50 UT and consists of two major sub-bursts (Bursts I and II), each lasting $\sim$20 minutes. The observations reveal three features: (1) Both burst components are highly intermittent, with well-defined upper and lower frequency cutoffs at $\sim$1–2 GHz and 200–300 MHz, respectively. (2) The maximum $T_{B}$ is $\sim$2$\times$1011 K for Burst I and $\sim$4$\times$1010 K for Burst II. (3) The polarization is left-handed with levels reaching $\sim$70–100%, with an overall increasing trend with frequency. Based on these observations, we conclude that the emission is coherent since the incoherent gyrosynchrotron radiation of solar flares cannot yield such intermittency with high $T_{B}$ and strong polarization at such frequencies (Dulk, 1985; Gary & Keller, 2004).
Fig. 2 and the animation 1 present the eruption recorded by AIA at 94, 131, and 171 Å, superposed by the NRH source contours at 8 frequencies from 228 to 445 MHz. This event was associated with an M7.1-class flare originating from NOAA Active Region (AR) 11302. The flare started at 12:33 UT, peaked at 13:17 UT, and ended at 14:10 UT. The AR has a complex multipolar topology with three sunspots. The event was associated with a halo coronal mass ejection. Both 94 and 131 Å data show the eruption of a hot structure and the rise of heated post-flare loops (Fu et al., 2023).
@@
Using the NLFFF extrapolation method, we deduced the coronal field lines of the AR (Figs. 2d and 2h), and overlaid them onto the HMI magnetogram. The field lines are color-coded to represent the section where the field strength varies from $\sim$40.7 G (blue) to $\sim$79.5 G (red), which correspond to the local second harmonic gyrofrequency (2 $\Omega_{ce}$) increasing from 228 to 445 MHz.
We draw two major conclusions: (1) NRH sources of both bursts (I and II) lie along the field lines pointing toward the sunspot with strong left-handed polarization (Fig. 1c), so both bursts are of X mode, and (2) the sources align well with the colored 2 $\Omega_{ce}$ section. Thus, the most likely radiation mechanism is the harmonic X mode via ECME since the alternative coherent plasma emission process would produce O mode for the fundamental branch or weak polarization for the harmonic branch (Dulk, 1985; Melrose et al., 1978; Melrose, 2017; Ni et al., 2020; Chen et al., 2022; Zhang et al., 2022).
4.2 Modeling particle transport for the 20110924 event
@@
To simulate the transport of energetic electrons that stream down toward the sunspot after being accelerated high in the corona, we release a population of Gaussian-distributed beam-like energetic electrons in the upper loop section (see the yellow sphere in Fig. 2h at an altitude of 160 Mm). The electrons, with the distribution of $f(\mu)\sim\exp\left(-\frac{(\mu-1)^{2}}{0.01}\right)$ where $\mu=\cos\theta$, first stream toward the negative sunspot and then get mirrored at different altitudes (Fig. 3b and the animation 2). As shown in Fig. 3c, a large fraction of 20–30 keV electrons get mirrored at heights above $\sim$10–20 Mm within 2 seconds after release, among them $\sim$85(97%) have pitch angles exceeding 90 (60). Eventually, $\sim$89.7of all injected electrons are mirrored at heights between 20 and 100 Mm (Fig. 3d). This result also holds for 10–20 keV, 30–50 keV, and 50–100 keV electrons since the mirroring point only depends on the initial pitch angles of particles.
As expected, the electrons within 20–100 Mm exhibit a shell-like VDF (Fig. 4), with significant positive gradient in the perpendicular direction. This drives the ECME to release both the fundamental X mode (X1) and the harmonic X mode (X2). According to the latest particle-in-cell simulation (Yousefzadeh et al., 2021) with similar numerical setup, energetic electrons with such shell-like VDFs excite the X2 mode more efficiently than the X1 mode. In addition, X1 experiences strong absorption at the corresponding second harmonic layer and thus cannot escape freely (Melrose & Dulk, 1982; Robinson, 1989; McKean et al., 1989; Melrose & Wheatland, 2016). Therefore, the X2 mode is relevant to the t-IVdm bursts observed from the Earth. Note that the ECME-radiating energetic electrons within such converging field are characterized by the shell or strip-like VDFs, rather than the generally-presumed loss-cone distribution.
@@
4.3 The 20120614 event
We further investigated the t-IVdm event on 14 June 2012 (see Appendix). The burst displays intermittent characteristics with a maximum $T_{B}$ of $\sim$2$\times$1010 K and exhibits strong levels of left-handed polarization (50–100%). During most time of the burst, the NRH sources are above a negative spot and coincide spatially with the corresponding extrapolated 2 $\Omega_{ce}$ section. This further supports the burst is excited via ECME at the X2 mode.
5 Discussion
@@
For the other 58 events collected here, we cannot determine the magnetic field orientation within the sources or the sources are characterized by weak polarizations. So their radiation mechanisms remain elusive. For events occurring after 2014, we cannot determine the sense of polarization because NRH stopped calibrating its polarization data since then. Further development of decimetric radio heliograph with well-calibrated data of high spatial and spectral resolution is demanded for a better understanding of such events.
The authors are grateful to the GOES, NRH, SDO/AIA, and HMI consortia for providing the data. We thank the Institute for Data Science FHNW Brugg/Windisch, Switzerland, for providing the e-Callisto data. SDO is a mission of NASA’s Living With a Star Program. This work was supported by the National Natural Science Foundation of China (12303061, 12127901, 12333009, 42074203 and 12203031), the National Key R&D Program of China under grant (2022YFF0503002 and 2022YFF0503000), the Shandong Natural Science Foundation of China (ZR2023QA074), and the China Postdoctoral Science Foundation (2023T160385, 2022M711931, and 2022TQ0189). A.K. was supported by the Ministry of Science and Higher Education of the Russian Federation. We acknowledge the computational resources provided by the Beijing Super Cloud Computing Center (BSCC, URL: http://www.blsc.cn).
The t-IVdm event on June 14 2012 Fig. S1 presents the spectra of the event, along with the NRH data of $T_{B}$ and polarization levels. The spectra span from $\sim$150 MHz to above 5 GHz, and can be divided into two components: the gyrosynchrotron emission at higher frequencies ($\sim$500 MHz – $>$5 GHz) peaking around 2 GHz, and the t-IVdm emission below $\sim$500 MHz. The t-IVdm burst starts from 13:40 to 15:20 UT, with $T_{B}$ decreasing with increasing frequency, being distinct from the usual optically thick gyrosynchrotron emission. According to Fig. S1, the t-IVdm burst exhibits the following characteristics: (1) It is intermittent with a well-defined lower frequency cutoff of $\sim$150–200 MHz, and the upper frequency cutoff is $\sim$500 MHz when it can be clearly separated from the gyrosynchrotron component. (2) The maximum $T_{B}$ is $\sim$2 $\times$ 1010 K. (3) At most NRH frequencies, the burst manifests strong left-handed polarization with a level of $\sim$50–100%. At 445 MHz from 14:14 to 14:16 UT, it presents a right-handed polarization with a level of $\sim$30%. Fig. S2, along with the animation 3, display the eruption observed by AIA at 94, 131, and 171 Å, superposed by the NRH source contours at 9 frequencies from 150 to 445 MHz. The event was associated with two M-class flares (M1.5 and M1.9) and a halo coronal mass ejection originating from NOAA AR 11504 that has two large spots with opposite polarities. The M1.5-class flare began at 12:47 UT, peaked at 14:08 UT, and ended at 15:00 UT while the M1.9-class flare started at 12:52 UT, reached its peak at 14:35 UT, and ended at 15:56 UT. Both 94 and 131 Å data reveal the eruption of hot structures, followed by heated post-flare loops. Similar to the event on 24 September 2011, the NRH sources align well with higher-frequency sources being closer to the disk. During the burst, the sources move systematically with at least two leaps. Before 13:50 UT, the sources lie above the east footpoint of the flaring loops, right above the left negative sunspot. Subsequently, the sources move to the region above the flaring loops. From $\sim$14:10 to $\sim$14:17 UT, they are located near the west footpoint of the loops, above the right positive sunspot. At $\sim$14:20 UT, the sources return to the region above the left spot. Figs. S2d and S2h present the extrapolated field lines and NRH sources, overplotted onto the HMI magnetogram. The field lines are color-coded to depict the section where the field strength varies from $\sim$26.8 G (blue) to $\sim$79.5 G (red). The corresponding second harmonic gyrofrequency (2 $\Omega_{ce}$) increases from 150 to 445 MHz. The NRH sources align well with the field lines pointing toward the sunspot and being co-spatial with the colored 2 $\Omega_{ce}$ section. As mentioned, the sources at 445 MHz show moderate level of right-handed polarization from 14:14 to 14:16 UT. This occurs when the sources are close to the positive spot, indicating the radiation is also in X mode. We summarize that the 20120614 burst is also in the X2 mode radiated coherently via ECME.

### v2

# A Flare-related Decimetric Type-IV Radio Burst Induced by the X2 Radiation of Electron Cyclotron Maser Emission

@@ Abstract
The radiation mechanism of decimetric wideband and pulsating radio bursts from the Sun (in terms of decimetric type-IV (t-IVdm) burst) and other flaring stars is a long-standing problem. Early investigations were based on the leading-spot hypothesis for the sun and yielded contradictory results. Here, we analyzed the flare-associated t-IVdm burst on 20110924 with medium-strong levels of polarization and from sources near a sunspot. We found that the emission is intermittent and the maximum $T_{B}$ exceeds 1011 K, with well-defined upper and lower frequency cutoffs. The radio sources are left-handed polarized, located above the sunspot with a negative polarity. The sources align well with the sites of the second harmonic of the local electron gyrofrequency. These findings provide essential evidence that the burst is induced by the electron cyclotron maser emission (ECME) in the harmonic X mode. We further modeled the transport of downward-streaming energetic electrons along a coronal loop and found most electrons get mirrored within the specific altitude range of 20–100 Mm. This explains why such bursts tend to have well-defined spectral ranges. We also found the ECME-radiating energetic electrons exhibit a shell-like VDF instead of the generally-presumed loss-cone distribution. The study greatly expands the application of ECME in solar radio astronomy and provides solar samples for similar bursts from other flaring stars.
@@ 1 Introduction
Due to the close proximity of the Sun to the Earth, solar radio astronomy has the advantage of observations with high spectral and spatial resolutions. This favors the determination of the underlying radiation mechanism. Even so, the exact origin of solar t-IVdm bursts and their substantial fine structures remain largely unresolved (Bernold, 1980; Benz, 1980; Slottje, 1982; Dulk, 1985; Güdel & Benz, 1988; Isliker & Benz, 1994; Fleishman et al., 2005; Vasanth et al., 2016; Morosan et al., 2019; White et al., 2024).
To determine the primary radiation mechanism, it is essential to identify the nature of the emission mode in terms of two freely escaping magnetoionic modes: extraordinary (X) or ordinary (O). This requires knowledge of the sense of polarization and orientation of the source magnetic field. Due to the lack of imaging data at decimetric wavelengths, to determine the source field orientation earlier studies (Aschwanden, 1986; Zlobec et al., 1987) were based on the leading sunspot hypothesis that assumes the bursts originate from the leading spots of the associated flares. Contradictory conclusions were drawn by these studies. As a result, distinct mechanisms have been proposed: plasma emission (PE; Kuijpers, 1974, 1980; Zaitsev & Stepanov, 1975; Benz & Kuijpers, 1976; Dulk, 1985), electron cyclotron maser emission (ECME; Benz & Tarnstrom, 1976; Aschwanden & Benz, 1988a, b), gyrosynchrotron radiation (e.g., Dulk 1973, 1985; Isliker & Benz 1994; Morosan et al. 2019), or resonant transition radiation (Nita et al., 2005; Fleishman et al., 2005)
Over the past eight decades of solar radio astronomy, only solar spikes that are highly transient (ms), highly-polarized, and extremly bright (with the brightness temperature ($T_{B}$) reaching above 1013 K) are widely attributed to ECME, while several other types of coherent solar bursts such as type-I, II, and III have been attributed to PE (e.g., Dulk 1985; Guedel & Zlobec 1991; Fleishman & Mel’nikov 1998; White et al. 2024). Note that some authors also suggested PE to be the origin of certain spike-like emission (e.g., Abalde et al. 2001; Chernov et al. 2001). Recent studies have reported a rare ECME-like event with week-long decimetric pulsations, which is not directly associated with any specific flare (Yu et al., 2024; Sharma et al., 2024). Here, we analyzed flare-induced t-IVdm bursts using multiwavelength data to identify the underlying radiation mechanism. In Section 2, we introduce how we selected the events, then we present the modeling methods of electron transport in Section 3. The results are presented in Section 4, and the conclusions and discussion are provided in Section 5.
@@ 2.1 Data and Instruments
The t-IVdm bursts were identified using the data from the 7m radio telescope of the Bleien observatory, which covers 175 to 870 MHz with a temporal resolution of 0.25 seconds. The Nançay Radioheliograph (NRH; Kerdraon & Delouis, 1997) offers imaging data with polarization measurements at 10 frequencies spanning from 150 to 445 MHz. The spatial resolution depends on the frequency and time of observation, being $\sim$2 at 445 MHz and $\sim$6 at 150 MHz in the summer, and up to three times larger along the north-south direction in the winter.
@@ 2.2 Event Selection
According to the general definition of t-IVdm bursts, we selected events with the maximum frequency exceeding 300 MHz, the bandwidth wider than 100 MHz, and the duration longer than 5 minutes. This yielded 60 events, from which we selected 37 flare-associated events (Table 1) according to the following criteria: the burst should be within 30 minutes around the peak of a flare, which originates from the active region (AR) associated with the NRH sources.
To identify the radiation mode, we only retained events with a high to moderate level of polarization ($>$30%) whose sources are located near a sunspot, with both latitude and longitude within 60, to ensure an unambiguous determination of the source-field orientation. The field orientation is further verified by the nonlinear force-free field (NLFFF) extrapolation. After applying these criteria, we identified 2 events that occurred on 20110924 and 20120614, respectively. For the 20120614 event, the burst passes through the complex loop system of the AR, so we cannot determine the emission mode unambiguously due to the possible reversal of polarization (see Section 5). Thus we only kept the 20110924 event here.
3 Methods of Modeling Electron Transport Along the Loop
@@ 3.1 Magnetic Field Processing and Extrapolation
We analyzed the HMI vector magnetic field data (“hmi.B_720s”), which cover the full solar disk with a temporal cadence of 720 s and a spatial resolution of 0.505′′ (Schou et al., 2012). The coronal magnetic field was reconstructed using the magneto-frictional method (Guo et al., 2016). To do this, we first corrected the projection effect (Gary & Hagyard, 1990) since the AR is away from the disk center, and then removed the residual force on the magnetograms using a preprocessing method (Wiegelmann et al., 2006) with optimized parameters, ensuring that the extrapolation boundary remained consistent with observations. We applied two metrics (Wheatland et al., 2000) to check whether the magnetic field reaches the force-free state or not. The obtained force-freeness metrics for both events are lower than 0.25, and the divergence-freeness metrics are close to 8 $\times$ 10-5 , similar to previous studies (Zhong et al., 2019).
3.2 Modeling of Electron Transport
@@ 4.1 Multiwavelength data analysis for the 20110924 event
Fig. 1 presents the spectrum of the t-IVdm burst, along with the NRH data for $T_{B}$ and polarization. The event occurred from 12:50 to 13:50 UT and consists of two major sub-bursts (Bursts I and II), each lasting $\sim$20 minutes. The observations reveal three features: (1) Both burst components are highly intermittent, with well-defined upper and lower frequency cutoffs at $\sim$1–2 GHz and 200–300 MHz, respectively. (2) The maximum $T_{B}$ is $\sim$2$\times$1011 K for Burst I and $\sim$4$\times$1010 K for Burst II. (3) The polarization is left-handed with levels reaching $\sim$70–100%, with an overall increasing trend with frequency. Based on these observations, we conclude that the emission is coherent since the incoherent gyrosynchrotron radiation of solar flares cannot yield such intermittency with high $T_{B}$ $>$1011 K and strong polarization at such frequencies (Dulk, 1985; Gary & Keller, 2004).
Fig. 2 and the accompanying animation present the eruption recorded by AIA at 94, 131, and 171 Å, superposed by the NRH source contours at 8 frequencies from 228 to 445 MHz. This event was associated with an M7.1-class flare originating from NOAA AR 11302. The flare started at 12:33 UT, peaked at 13:17 UT, and ended at 14:10 UT. The AR has a complex multipolar topology with three sunspots. The event was associated with a halo coronal mass ejection. Both 94 and 131 Å data show the eruption of a hot structure and the rise of heated post-flare loops (Fu et al., 2023).
@@
Using the NLFFF extrapolation method, we deduced the coronal field lines of the AR (Figs. 2d and 2h), and overlaid them onto the HMI magnetogram. The field lines are color-coded to represent the section where the field strength varies from $\sim$40.7 G (blue) to $\sim$79.5 G (red), the corresponding harmonic gyrofrequency (2 $\Omega_{ce}$) increases from 228 to 445 MHz.
We draw two major conclusions: (1) NRH sources of both bursts (I and II) lie along the field lines pointing toward the sunspot with strong left-handed polarization (Fig. 1c), so both bursts are of X mode, and (2) the sources align well with the colored 2 $\Omega_{ce}$ section. Thus, the most likely radiation mechanism is the harmonic X mode (X2) via ECME since the alternative coherent plasma emission process would produce O mode for the fundamental branch or weak polarization for the harmonic branch (Dulk, 1985; Melrose et al., 1978; Melrose, 2017; Ni et al., 2020; Chen et al., 2022; Zhang et al., 2022).
4.2 Possibility of Polarization Reverse for the 20110924 event
According to the theory of mode coupling, the polarization sense of the radiation will be reversed when it propagates through a weak-coupling quasi-transverse (QT) layer, i.e., the radiation frequency is lower than the transition frequency of the QT layer (Cohen, 1960; Benz et al., 1977; White et al., 1992).
To determine whether the polarization of the burst is reversed by the propagation effect, we checked the AR polarization before and after the burst (Fig. 3). The polarization in the burst source is generally left-handed, with the same sense to the burst. The non-eruptive AR radiates in radio via gyroresonance and bremsstrahlung of thermal electrons, mainly in the X mode (see, e.g., Dulk 1985; White & Kundu 1997; Gary & Keller 2004).
We further examined the height of the ejecta during the burst. According to the Large Angle and Spectrometric Coronagraph (LASCO C2; Brueckner et al. 1995), the coronal mass ejection as a whole has moved beyond 2 $R_{s}$ above the solar surface during the burst. This indicates the complex eruptive structure has moved beyond the weak-coupling regime that is roughly below 0.9 $R_{s}$ for the current event. This deduction agrees with AIA data of the eruption. So during the burst, the complex eruptive structure cannot result in the sense reversal of polarization. Combining the above analyses, we suggest that the burst radiation is also in X mode.
4.3 Modeling electron transport for the 20110924 event
@@
We assume the electrons are accelerated by the flare reconnection that takes place relatively high in the corona. The electrons then flow downward and upward. The Hard X-rays of solar flares and some drifting-toward-higher-frequency type-III bursts are evidence of these downward streaming energetic electrons (e.g., Aschwanden et al. 1995; Reid & Ratcliffe 2014). To simulate the transport of these energetic electrons, we release a population of Gaussian-distributed beam-like energetic electrons in the upper loop section (see the yellow sphere in Fig. 2h at an altitude of 160 Mm). The electrons, with the distribution of $f(\mu)\sim\exp\left(-\frac{(\mu-1)^{2}}{0.01}\right)$ where $\mu=\cos\theta$, first stream toward the negative sunspot and then get mirrored at different altitudes (Fig. 4b and the animation 2). As shown in Fig. 4c, a large fraction of 20–30 keV electrons get mirrored at heights above $\sim$10–20 Mm within 2 seconds after release, among them $\sim$85(97%) have pitch angles exceeding 90 (60). Eventually, $\sim$89.7of all injected electrons are mirrored at heights between 20 and 100 Mm (Fig. 4d). This result also holds for 10–20 keV, 30–50 keV, and 50–100 keV electrons since the mirroring point only depends on the initial pitch angles of electrons due to the neglect of wave/turbulence-particle interaction. In the real corona, the mirroring point depends on such interaction and other factors, such as variation of the magnetic field, Coulomb collisions, etc.
As expected, the electrons within 20–100 Mm exhibit a shell-like VDF (Fig. 5), with significant positive gradient in the perpendicular direction. This type of VDF can drive the ECME to release both the fundamental O or X mode (O1 or X1) and X2. According to the latest particle-in-cell simulation (Ning et al., 2021; Yousefzadeh et al., 2021, 2022) with similar numerical setup, energetic electrons with such shell-like VDFs excite the X2 mode much more efficiently than the fundamental modes that will experience strong damping at the corresponding second harmonic absorption layer, while the absorption of X2 mode is much weaker (Melrose & Dulk, 1982; Robinson, 1989; McKean et al., 1989; Melrose & Wheatland, 2016). This agrees with the above suggestion that the t-IVdm radiation is the X2 mode excited via the ECME process.
Note that the VDF of the ECME-radiation electrons is shell-like, rather than being the loss-cone type, as presumed by most earlier studies(e.g., Winglee & Dulk 1986; Aschwanden & Benz 1988a; Stupp 2000). According to our simulation, the shell-like distribution is mainly caused by the mirroring motion of energetic electrons along the converging sunspot field. We point out that the exact details of the formation of particle VDF depend on the conditions of wave-particle interaction along the loop, the shape of the initial VDF, and other factors such as the time and location of the injection.
@@
5 Conclusions and Discussion
Here, we investigated the radiation mechanism of flare-induced t-IVdm bursts using multi-wavelength data. We analyzed the burst on 20110924 with medium-strong levels of polarization and originating from sources near a sunspot. The emission is intermittent and the maximum $T_{B}$ exceeds 1011 K, with well-defined upper and lower frequency cutoffs at $\sim$1–2 GHz and 200–300 MHz, respectively. The radio sources of the burst show left-handed polarization, and lie along the field lines pointing toward the sunspot with a negative polarity. According to the extrapolation of the coronal magnetic field, we found a good correlation between the sources and the sites of the second harmonic of the local electron gyrofrequency. These findings provide substantial evidence that the burst is the X2 radiation of ECME.
We further modeled the electron transport along the coronal loop. We found that most electrons get mirrored in converging sunspot fields within the altitude range of 20–100 Mm. This agrees with the well-defined spectral ranges of such bursts. The ECME-radiating energetic electrons exhibit a shell-like VDF, rather than the generally-presumed loss-cone distribution.
[…]
According to one earlier analysis (Zlotnik, 1981), the harmonic plasma emission could be in X mode if its mother mode, i.e., the Langmuir waves, propagates isotropically, and the polarization degree never exceeds 40%. Yet with most updated particle-in-cell simulations (e.g., Chen et al. 2022; Zhang et al. 2022), the Langmuir waves driven by electron beams propagate mostly quasi-parallel to the beam. So the above result does not hold for most beam-driven situation. So we have excluded the possibility of harmonic plasma emission to account for the observed highly-polarized t-IVdm burst.
@@ 5 Conclusions and Discussion
For the other 59 events collected here, we cannot determine the magnetic field orientation within the sources or the sources are characterized by weak polarizations. So their radiation mechanisms remain elusive. For events occurring after 2014, we cannot determine the sense of polarization because NRH stopped calibrating its polarization data since then. Further development of decimetric radio heliograph with well-calibrated data of high spatial and spectral resolution is demanded for a better understanding of such events.
The authors are grateful to the GOES, NRH, SDO/AIA, and HMI consortia for providing the data. We thank the Institute for Data Science FHNW Brugg/Windisch, Switzerland, for providing the e-Callisto data. SDO is a mission of NASA’s Living With a Star Program. This work was supported by the National Natural Science Foundation of China (12303061, 12127901, 12333009, and 12203031), the National Key R&D Program of China (2022YFF0503002 and 2022YFF0503000), the Shandong Natural Science Foundation of China (ZR2023QA074), and the China Postdoctoral Science Foundation (2023T160385, 2022M711931, and 2022TQ0189). A.K. was supported by the Ministry of Science and Higher Education of the Russian Federation. We acknowledge the computational resources provided by the Beijing Super Cloud Computing Center (BSCC, URL: http://www.blsc.cn).